# Training on the Test Task

### v2

@@ 6 Related work
Data contamination or test-set contamination refers to any overlap between the training and the test data such that test results overestimate a model’s generalization performance. The scale and often little curation of present-day pretraining corpora exacerbates data contamination concerns in language model evaluations (Jiang et al., 2024). Consequently, data contamination is usually discussed in the technical reports accompanying model releases (Radford et al., 2019; Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023b). However, detecting and preventing data contamination is currently an open problem (Gunasekar et al., 2023; Yang et al., 2023b; Golchin and Surdeanu, 2023). Roberts et al. (2023) and Li and Flanigan (2024) find that models often perform better on datasets that were publicly available during model training. While almost all models that we consider were released after MMLU and GSM8K, we nonetheless find that, controlling for compute, more recent models perform better. These performance gains are unlikely to be driven solely by test set leakage and require additional explanation. In Section 3.2, we find evidence that that training on the test task may be a more dominant factor in benchmark performance than data contamination. A key insight of our work is that models can train on the test task (e.g., multiple-choice question answering) and do much better at a given benchmark (e.g., MMLU) without necessarily seeing any benchmark data. This defies traditionally notions of data contamination, which strictly refer to the leakage of benchmark data (Magar and Schwartz, 2022; Sainz et al., 2023; Dong et al., 2024). Moreover, whereas the core concern with data leakage is that benchmarks may inadvertently find their way to the training data, training on the test task is often a deliberate design choice (e.g., pre-training on instruction data).
@@
The effectiveness of fine-tuning on the training set accompanying LLM benchmarks is well-known (Wei et al., 2022a; Wang et al., 2022; Chung et al., 2024). Consequently, many influential instruction-tuning datasets contain or are partly derived from benchmark train data (Wei et al., 2022a; Honovich et al., 2022; Mukherjee et al., 2023). Li and Flanigan (2024) identify small amounts of benchmark-specific data in the publicly available Alpaca (Taori et al., 2023) and Vicuna (Chiang et al., 2023) instruction-tuning sets. Zhou et al. (2023b) empirically analyze the effects of fine-tuning on benchmark-specific data and warn about its impacts on benchmark validity. To circumvent these issues, recent work has focused on indirect indicators of broader data contamination, such as a lack of robustness to task transformations (Wu et al., 2023), or underperformance on benchmarks with novel task combinations (Yu et al., 2023a). In contrast, we find evidence for training on the test task without the need for explicitly identifying specific data points used at training time, or modifying tasks. In addition, our proposed method of fine-tuning on task data prior to evaluation allows us to quantify and correct for the effects of training on the test task on benchmark performance.
@@
Emergent capabilities (Wei et al., 2022b; Ganguli et al., 2022) refer to levels of model performance at large scales that cannot be easily predicted by extrapolating from smaller scales. Wei et al. (2022b) report emergent capabilities for various benchmarks including MMLU and GSM8K (Srivastava et al., 2022). However, Srivastava et al. (2022); Schaeffer et al. (2024b) find that the log-probability of the correct answer often improves smoothly, even when other metrics seem to show emergence. Lu et al. (2023) argue that most emergent capabilities can be explained by in-context-learning. Schaeffer et al. (2024a) argue that emergent capabilities are mostly an artifact of non-linear and discontinuous evaluation metrics like accuracy. In contrast, we find signs of emergence on MMLU even when using continuous metrics like the Brier score. Similarly to our findings, Snell et al. (2024) show that increasingly fine-tuning on the test task shifts the point of emergence to smaller compute scales.
@@ 7 Discussion
The 1968 Olympics took place in Mexico City at the significant altitude of 2340 meters, higher than Australia’s tallest peak. Runners who had trained at altitude in their home countries were better prepared to compete in Mexico City’s conditions, as it turned out. But the hotly debated results of the Games did not lead the organizers to prohibit training at natural altitude. Instead, they let everyone do it; and athletes came to consider altitude training an excellent way to train.
@@
Our work demonstrates that comparisons of different models are confounded by the choice of training data and training practices. Different model families vary in the degree that they were—implicitly or explicitly—trained on various test tasks. A relatively small amount of task data can have a disproportional large effect on benchmark performance. It therefore makes little sense to compare model performance at face value without accounting for how the training data relate to the test task.
Training on the test task also has profound implications for the study of emergent behavior. After training on the test task, model capabilities become predictable at smaller scales. This greatly reduces the unpredictability associated with emergence, notably without any change in the metric.
@@ A.3 Robustness check on the hyperparameter search budget
A potential concern is that older models may appear to match the performance of newer models because the fine-tuning hyperparameters could be systematically more favorable to the older models. To address this concern, we recreate our main experiment by fine-tuning older models with their original learning rates while selecting for the newer models the optimal learning rate in the sweep that leads to highest MMLU performance. That is, we deliberately give newer models a systematic advantage in terms of hyperparameter search budget. We plot the results in Figure 9. The estimated effect size of model recency on benchmark performance shifts from $\widehat{\theta}=-0.005$ to $\widehat{\theta}=0.005$. Therefore, despite giving newer models a systemic advantage in terms of hyperparameter search budget, the effect size remains both remains small and not statistically significant.
@@ Appendix B Causal interpretation of our findings
The key obstacle to our analysis is that test task training $T$ is unobservable. Firstly, because practitioners are typically not transparent about their designs choices, including the pretraining data. Secondly, because the extent to which different training practices might amount to test task training is unclear. Nonetheless, by fine-tuning on task-specific data we are able to intervene on the extent to which models train on the test task.
@@
We model fine-tuning as a hard intervention $\textrm{do}(T=t)$ (Pearl, 2009). The specific magnitude of the intervention $t$ need not be quantified. Instead, the key assumption is that by fine-tuning on the same, sufficient amount of task data, all models will have received the same amount of test task training. Since some base models may have already trained on the test task prior to fine-tuning, this assumption only holds if test task training saturates and we train on enough task data to reach saturation. The fact that our task-specific datasets allow older models to match the performance of newer models provides some evidence that we train on enough task-specific data to reach saturation.
We draw inspiration from scaling laws (Kaplan et al., 2020) and model relationship between pretraining compute and its causal descendants as pice-wise log-linear:
@@
where $\alpha,\psi,\eta^{\prime}$ are the fit’s parameters and $\epsilon$ is random noise. The coefficient $\psi$ corresponds to the direct effect $N\rightarrow A$ of model recency on benchmark accuracy. We additionally regress on the difference in accuracy pre and post intervention
@@ C.1 Adjusting the temporal cutoff by a few months
We repeat the analysis of Section 2 for two additional temporal splits: September 2023 and January 2024, and present the results in Figure 11 and Figure 12, respectively. Our results are robust to shifting the temporal cutoff by a few months. That is, our findings indicate that practitioners started adopting design choices around late 2023 that amount to models training on the test task much more, which is consistent with models’ technical reports starting to mention the use of benchmark or instruction data at pre-training time. Choosing specifically the month of November as the cut-off is therefore not critical for our analysis.
@@ C.3 How similar are newer models to older, fine-tuned models?
In Section 3.1 we fine-tune older models on the test task and we demonstrate that the differences in benchmark performance between the fine-tuned and non fine-tuned models resemble those between newer and older models. In this section we provide further evidence that newer models resemble older, fine-tuned models.
@@
We perform a quantitative analysis consisting in discriminating between the older models and the newer models based on their pretraining compute and benchmark accuracy. That is, we construct a tabular dataset where rows are models and columns are their corresponding pretraining compute, benchmark accuracy, and whether the model was trained after November 2023. We then train a classifier aiming to predict model recency from compute and accuracy. Intuitively, if the performance of older models is very different form that of newer models, then we would obtain high prediction accuracy (i.e., the two classes are highly separable). Note that prediction accuracy provides a lower bound on the total variation (TV) distance between the distributions of compute and accuracy of older and newer models.
@@ Appendix D Results for the OpenLLM Leaderboard v2
The creators of the OpenLLM Leaderboard cite contamination as a key motivation for releasing the v2 revision. They note that a key criteria in choosing the benchmarks of the HF leaderboard v2 was lack of contamination in models as of today. In particular, Fourrier et al. (2024b) claim that current models are not contaminated for GPQA, MuSR, and MMLU Pro: GPQA due to the gating of the test set, and MuSR and MMLU Pro due to their “youth”. Fourrier et al. (2024b) succinctly express their concern as regards to data contamination in the HF leaderboard v1:
@@
We evaluate all models on MMLU Pro, GPQA, BBH, MuSR and MATH Lvl 5. We use the LM Evaluation Harness library in identical fashion to the HF leaderboard v2. We do not evaluated on IFEval since it tests for instruction following and we evaluate base models. We additionally evaluate the models that we fine-tuned in Section 2.1 for multiple choice question answering and mathematical reasoning. This gives us models’ adjusted benchmark scores after training on multiple choice question answering and mathematical reasoning. For MATH Lvl 5, we use the models fine-tuned on mathematical data, whereas for MMLU Pro, GPQA, BBH and MuSR we use the models fine-tuned on multiple choice question answering. The fine-tuning datasets were not adapted to the new benchmarks in the HF leaderboard v2, thus giving a valuable insight into how well these task-relevant datasets generalize beyond MMLU and GSM8K.
We plot in Figure 15 models benchmark scores pre and post post adjustment. We find that newer models significantly outperform older ones in all five benchmarks after controlling for pretraining compute. The differences in performance are smaller in absolute terms than those measured for MMLU (0.073) and GSM8K (0.191). This is in part because these benchmarks are “harder”, meaning also smaller differences in performance between the best and worst model. For this reason, we also report the difference between newer and older models relative to the difference between the best and worst model. This relative difference is 13.7% for MMLU Pro, 14.5% for GPQA, 12.1% for MuSR, 9.7% for BBH, and 10.0% for MATH Lvl 5, compared to 15.3% for MMLU and 25.0% for GSM8K. Therefore, newer models overperform in MMLU Pro, GPQA and MuSR about as much as they do for MMLU, and somewhat less for BBH and MATH Lvl 5.
@@
One possible reason for the fact that the adjustment for MMLU Pro and MATH Lvl 5 is not as effective as for MMLU and GSM8K is that the fine-tuning examples are simply not as relevant for MMLU Pro and MATH Lvl 5. For example, the questions and answers in MATH Lvl 5 contain much more LaTeX equation formatting than our mathematical reasoning fine-tuning dataset. Similarly, our multiple choice fine-tuning dataset contains mostly questions with 4 answer choices, whereas all MMLU Pro questions have 10 answer choices. Thus, models are primarily fine-tuned to answer “A”, “B”, “C”, and “D” but not “E”, “F”, “G”. We modify MMLU Pro to contain questions with 4 answer choices by randomly discarding 6 of the incorrect answer choices. We evaluate models pre and post adjustment and plot the results in Figure 16. We observe that the difference in performance between newer and older models after adjustment reduces from 0.024 to 0.016, and is no longer statistically significant. This observation suggests that fine-tuning one more relevant task-data might further reduce the gap between newer and older models in MMLU Pro and MATH Lvl 5.
@@ Discussion.
Fourrier et al. (2024b) cite newer models overperforming in the HF leaderboard v1 due to being “possibly trained on benchmark data or on data very similar to benchmark data” as a major reason for the HF leaderboard v2 revision. We however find evidence that training on the test task is also a confounder for the newly included benchmarks. Specifically, the difference in performance between newer and older models is significant for MMLU Pro, GPQA, MuSR, BBH and MATH Lvl 5, and these differences reduce after adjusting by fine-tuning on the test task.
Fourrier et al. (2024b) explicitly highlight GPQA and MuSR as benchmarks likely unaffected by contamination, the former due to being gated and latter due to its “youth”. Not only do newer models significantly outperform older ones in GPQA and MuSR, but these differences in performance fully vanish after fine-tuning on the test task. That is, newer models likely overperform in GPQA and MuSR precisely due to training on the test task.
@@
Appendix E Additional figures on emergence
@@ Appendix F Results for instruction-tuned and chat models
We evaluate 36 instruct and chat models, see Appendix A.1.1. Our findings for base models presented in the main text generalize remarkably well to instruction-tuned and chat models, see Figure 19. Newer instruct/chat models substantially outperform older instruct/chat models. However, after fine-tuning all models on the same amount of task-specific data, performance between newer and older instruct/chat models equalizes.

### v3

@@ 6 Related work
Data contamination or test-set contamination refers to any overlap between the training and the test data such that test results overestimate a model’s generalization performance. The scale and often little curation of present-day pretraining corpora exacerbates data contamination concerns in language model evaluations (Jiang et al., 2024). Consequently, data contamination is usually discussed in the technical reports accompanying model releases (Radford et al., 2019; Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023b). However, detecting and preventing data contamination is currently an open problem (Gunasekar et al., 2023; Yang et al., 2023b; Golchin and Surdeanu, 2023). Roberts et al. (2023) and Li and Flanigan (2024) find that models often perform better on datasets that were publicly available during model training. While almost all models that we consider were released after MMLU and GSM8K, we nonetheless find that, controlling for compute, more recent models perform better. These performance gains are unlikely to be driven solely by test set leakage and require additional explanation. In Section 3.2, we find evidence that that training on the test task may be a more dominant factor in benchmark performance than data contamination. A key insight of our work is that models can train on the test task (e.g., multiple-choice question answering) and do much better at a given benchmark (e.g., MMLU) without necessarily seeing any benchmark data. This defies traditional notions of data contamination, which strictly refer to the leakage of benchmark data (Magar and Schwartz, 2022; Sainz et al., 2023; Dong et al., 2024). Moreover, whereas the core concern with data leakage is that benchmarks may inadvertently find their way to the training data, training on the test task is often a deliberate design choice (e.g., pre-training on instruction data).
@@
The effectiveness of fine-tuning on the training set accompanying LLM benchmarks is well-known (Wei et al., 2022a; Wang et al., 2022; Chung et al., 2024). Consequently, many influential instruction-tuning datasets contain or are partly derived from benchmark train data (Wei et al., 2022a; Honovich et al., 2022; Mukherjee et al., 2023). Li and Flanigan (2024) identify small amounts of benchmark-specific data in the publicly available Alpaca (Taori et al., 2023) and Vicuna (Chiang et al., 2023) instruction-tuning sets. Zhou et al. (2023b) empirically analyze the effects of fine-tuning on benchmark-specific data and warn about its impacts on benchmark validity. To circumvent these issues, recent work has focused on indirect indicators of broader data contamination, such as a lack of robustness to task transformations (Wu et al., 2023), or underperformance on benchmarks with novel task combinations (Yu et al., 2023a). In contrast, we find evidence for training on the test task without the need for explicitly identifying specific data points used at training time, or modifying tasks. In addition, our proposed method of fine-tuning on task data before evaluation allows us to quantify and correct for the effects of training on the test task on benchmark performance.
@@
Emergent capabilities (Wei et al., 2022b; Ganguli et al., 2022) refer to levels of model performance at large scales that cannot be easily predicted by extrapolating from smaller scales. Wei et al. (2022b) report emergent capabilities for various benchmarks including MMLU and GSM8K (Srivastava et al., 2022). However, Srivastava et al. (2022); Schaeffer et al. (2024b) find that the log-probability of the correct answer often improves smoothly, even when other metrics seem to show emergence. Rogers and Luccioni (2024) question the dominant definition of emergence and emphasize the importance of relating the training data to the test data before making claims about emergence. Lu et al. (2023) argue that most emergent capabilities can be explained by in-context-learning. Gadre et al. (2024) find that a model’s perplexity on its pre-training data reliably predicts its average downstream performance. However, their analysis does not include the MMLU and GSM8K benchmarks, which are central to our work. Schaeffer et al. (2024a) argue that emergent capabilities are mostly an artifact of non-linear and discontinuous evaluation metrics like accuracy. In contrast, we find signs of emergence on MMLU even when using continuous metrics like the Brier score. Similarly to our findings, Snell et al. (2024) show that increasingly fine-tuning on the test task shifts the point of emergence to smaller compute scales.
@@ 7 Discussion
The 1968 Olympics took place in Mexico City at the significant altitude of 2340 meters, higher than Australia’s tallest peak. Runners who had trained at altitude in their home countries were better prepared to compete in Mexico City’s conditions, as it turned out. But the hotly debated results of the Games did not lead the organizers to prohibit training at natural altitude. Instead, they let everyone do it, and athletes came to consider altitude training an excellent way to train.
@@
Our work demonstrates that comparisons of different models are confounded by the choice of training data and training practices. Different model families vary in the degree that they were—implicitly or explicitly—trained on various test tasks. A relatively small amount of task data can have a disproportionally large effect on benchmark performance. It therefore makes little sense to compare model performance at face value without accounting for how the training data relate to the test task.
Training on the test task also has profound implications for the study of emergent behavior. After training on the test task, model capabilities become predictable at smaller scales. This greatly reduces the unpredictability associated with emergence, notably without any change in the metric, thus largely disarming the ominous nature of emergence.
@@ A.3 Robustness check on the hyperparameter search budget
A potential concern is that older models may appear to match the performance of newer models because the fine-tuning hyperparameters could be systematically more favorable to the older models. To address this concern, we recreate our main experiment by fine-tuning older models with their original learning rates, while selecting for the newer models the optimal learning rate in the sweep that leads to highest MMLU performance. That is, we deliberately give newer models a systematic advantage in terms of hyperparameter search budget. We plot the results in Figure 10. The estimated effect size of model recency on benchmark performance shifts from $\widehat{\theta}=-0.005$ to $\widehat{\theta}=0.005$. Therefore, despite giving newer models a systemic advantage in terms of hyperparameter search budget, the effect size remains both small and not statistically significant.
@@ Appendix B Causal interpretation of our findings
The key obstacle to our analysis is that test task training $T$ is unobservable. Firstly, because practitioners are typically not transparent about their design choices, including the pretraining data. Secondly, because the extent to which different training practices might amount to test task training is unclear. Nonetheless, by fine-tuning on task-specific data, we can intervene on the extent to which models train on the test task.
@@
We model fine-tuning as a hard intervention $\textrm{do}(T=t)$ (Pearl, 2009). The specific magnitude of the intervention $t$ need not be quantified. Instead, the key assumption is that by fine-tuning on the same, sufficient amount of task data, all models will have received the same amount of test task training. Since some base models may have already trained on the test task before fine-tuning, this assumption only holds if test task training saturates, and we train on enough task data to reach saturation. The fact that our task-specific datasets allow older models to match the performance of newer models provides some evidence that we train on enough task-specific data to reach saturation.
We draw inspiration from scaling laws (Kaplan et al., 2020) and model the relationship between pretraining compute and its causal descendants as piecewise log-linear:
@@
where $\alpha,\psi,\eta^{\prime}$ are the fit’s parameters and $\epsilon$ is random noise. The coefficient $\psi$ corresponds to the direct effect $N\rightarrow A$ of model recency on benchmark accuracy. We additionally regress on the difference in accuracy pre- and post-intervention
@@ C.1 Adjusting the temporal cutoff by a few months
We repeat the analysis of Section 2 for two additional temporal splits: September 2023 and January 2024, and present the results in Figure 12 and Figure 13, respectively. Our results are robust to shifting the temporal cutoff by a few months. That is, our findings indicate that practitioners started adopting design choices around late 2023 that amount to models training on the test task more, which is consistent with models’ technical reports starting to mention the use of benchmark or instruction data at pre-training time. Choosing specifically the month of November as the cut-off is therefore not critical for our analysis.
@@ C.3 How similar are newer models to older, fine-tuned models?
In Section 3.1 we fine-tune older models on the test task, and we demonstrate that the differences in benchmark performance between the fine-tuned and non fine-tuned models resemble those between newer and older models. In this section we provide further evidence that newer models resemble older, fine-tuned models.
@@
We perform a quantitative analysis consisting in discriminating between the older models and the newer models based on their pretraining compute and benchmark accuracy. That is, we construct a tabular dataset where rows are models and columns are their corresponding pretraining compute, benchmark accuracy, and whether the model was trained after November 2023. We then train a classifier aiming to predict model recency from compute and accuracy. Intuitively, if the performance of older models is very different from that of newer models, then we would obtain high prediction accuracy (i.e., the two classes are highly separable). Note that prediction accuracy provides a lower bound on the total variation (TV) distance between the distributions of compute and accuracy of older and newer models.
@@ Appendix D Results for the OpenLLM Leaderboard v2
The creators of the OpenLLM Leaderboard cite contamination as a key motivation for releasing the v2 revision. They note that a key criterion in choosing the benchmarks of the HF leaderboard v2 was lack of contamination in models as of today. In particular, Fourrier et al. (2024b) claim that current models are not contaminated for GPQA, MuSR, and MMLU Pro: GPQA due to the gating of the test set, and MuSR and MMLU Pro due to their “youth”. Fourrier et al. (2024b) succinctly express their concern as regards to data contamination in the HF leaderboard v1:
@@
We evaluate all models on MMLU Pro, GPQA, BBH, MuSR and MATH Lvl 5. We use the LM Evaluation Harness library in an identical fashion to the HF leaderboard v2. We do not evaluate on IFEval since it tests for instruction following and we evaluate base models. We additionally evaluate the models that we fine-tuned in Section 2.1 for multiple choice question answering and mathematical reasoning. This gives us models’ adjusted benchmark scores after training on multiple choice question answering and mathematical reasoning. For MATH Lvl 5, we use the models fine-tuned on mathematical data, whereas for MMLU Pro, GPQA, BBH and MuSR we use the models fine-tuned on multiple choice question answering. The fine-tuning datasets were not adapted to the new benchmarks in the HF leaderboard v2, thus giving a valuable insight into how well these task-relevant datasets generalize beyond MMLU and GSM8K.
We plot in Figure 16 models benchmark scores pre- and post-post adjustment. We find that newer models significantly outperform older ones in all five benchmarks after controlling for pretraining compute. The differences in performance are smaller in absolute terms than those measured for MMLU (0.073) and GSM8K (0.191). This is in part because these benchmarks are “harder”, meaning also smaller differences in performance between the best and worst model. For this reason, we also report the difference between newer and older models relative to the difference between the best and worst model. This relative difference is 13.7% for MMLU Pro, 14.5% for GPQA, 12.1% for MuSR, 9.7% for BBH, and 10.0% for MATH Lvl 5, compared to 15.3% for MMLU and 25.0% for GSM8K. Therefore, newer models overperform in MMLU Pro, GPQA and MuSR about as much as they do for MMLU, and somewhat less for BBH and MATH Lvl 5.
@@
One possible reason for the fact that the adjustment for MMLU Pro and MATH Lvl 5 is not as effective as for MMLU and GSM8K is that the fine-tuning examples are simply not as relevant for MMLU Pro and MATH Lvl 5. For example, the questions and answers in MATH Lvl 5 contain much more LaTeX equation formatting than our mathematical reasoning fine-tuning dataset. Similarly, our multiple choice fine-tuning dataset contains mostly questions with 4 answer choices, whereas all MMLU Pro questions have 10 answer choices. Thus, models are primarily fine-tuned to answer “A”, “B”, “C”, and “D” but not “E”, “F”, “G”. We modify MMLU Pro to contain questions with 4 answer choices by randomly discarding 6 of the incorrect answer choices. We evaluate models pre- and post-adjustment and plot the results in Figure 17. We observe that the difference in performance between newer and older models after adjustment reduces from 0.024 to 0.016, and is no longer statistically significant. This observation suggests that fine-tuning one more relevant task-data might further reduce the gap between newer and older models in MMLU Pro and MATH Lvl 5.
@@ Discussion.
Fourrier et al. (2024b) cite newer models overperforming in the HF leaderboard v1 due to being “possibly trained on benchmark data or on data very similar to benchmark data” as a major reason for the HF leaderboard v2 revision. We, however, find evidence that training on the test task is also a confounder for the newly included benchmarks. Specifically, the difference in performance between newer and older models is significant for MMLU Pro, GPQA, MuSR, BBH and MATH Lvl 5, and these differences reduce after adjusting by fine-tuning on the test task.
Fourrier et al. (2024b) explicitly highlight GPQA and MuSR as benchmarks likely unaffected by contamination, the former due to being gated and the latter due to its “youth”. Not only do newer models significantly outperform older ones in GPQA and MuSR, but these differences in performance fully vanish after fine-tuning on the test task. That is, newer models likely overperform in GPQA and MuSR precisely due to training on the test task.
@@
Appendix E Additional figures
Figure 18: Training on the test task significantly alters model rankings on MMLU.
MMLU rankings
Training on the test task significantly alters model rankings on MMLU, with an average shift of 4.8 ranks and a maximum shift of 16 ranks.
@@ Appendix F Results for instruction-tuned and chat models
We evaluated 36 instruct and chat models, see Appendix A.1.1. Our findings for base models presented in the main text generalize remarkably well to instruction-tuned and chat models, see Figure 21. Newer instruct/chat models substantially outperform older instruct/chat models. However, after fine-tuning all models on the same amount of task-specific data, performance between newer and older instruct/chat models equalizes.